# A Survey on Error-Bounded Lossy Compression for Scientific Datasets

### v1

@@ 4. General-Purpose Error-Bounded Lossy Compressors
SZ (Di and Cappello, 2016; Tao et al., 2017c; Zhao et al., 2021; Liang et al., 2021; Di and Cappello, 2018) is a prediction-based error-bounded lossy compressor. In fact, it is not only a compression library/software but also a flexible composable framework allowing users to customize specific compression pipelines according to their datasets or use cases. SZ’s compression pipeline generally is composed of four stages: pointwise data prediction, quantization, variable-length encoding and lossless encoding. For different domain datasets and use cases, the SZ developers have developed many predictors, including Lorenzo, linear regression, and dynamic spline interpolation (Zhao et al., 2021), using ZFP transform as a predictor (Liang et al., 2019a), as listed in Section 3.1.
In addition to the data prediction stage, SZ developers explored the possibilities of improving compression capability by other stages, for example, different quantization methods such that the users can set various error bounds for multiple value ranges in one dataset (Liu et al., 2021b, c, 2022a). SZ adopts Huffman+Zstd (Di and Cappello, 2016; Tao et al., 2017c; Liang et al., 2018b) to compress the quantization bins because this method projects the best trade-off between the compression ratio and compression speed (Liu et al., 2021d). For pointwise relative error bound compression in SZ (Liang et al., 2018a), SZ performs a preprocessing step to transform the original data domain to the logarithm domain and then execute absolute error-bounded compression on top of it. However, the logarithm may significantly delay the overall compression/decompression performance, which was fixed by an efficient fusion of logarithm and quantization thereafter (Zou et al., 2019, 2020). Since SZ adopts a fixed number of quantization bins for each compression, some data points may be outliers (i.e., outside the quantization range). These outliers are called ‘unpredictable data’, which will be compressed in a separate way (e.g., using bit-wise truncation).
4.2. ZFP
@@
ZFP (Lindstrom, 2014) is a transform-based error-bounded lossy compressor, which supports two error control methods: fixed-accuracy (i.e., absolute error bound) and fixed precision. ZFP splits the whole dataset into many fixed-size blocks (e.g., 4$\times$4$\times$4 for a 3D dataset) and then executes three steps in each block: (1) preprocessing (PDT): align the values in a block to a common exponent and convert the floating-point values to a fixed-point representation; (2) (near)orthogonal block transform (OWT): use orthogonal transform to decorrelate data; and (3) embedded coding (BPC): order and encode the transform coefficients by the embedded coding. To achieve the best trade-off between decorrelation efficiency and speed, developers of ZFP explored multiple transforms using a parametric description and identified a near-orthogonal one to use in practice. Their embedded encoding is a variation of BPC, where the coefficients are divided into separate groups based on their locations and then encoded in the granularity of a group. In general, ZFP features high compression and decompression performance on both CPUs and GPUs because of the performance optimization strategies in its implementation, such as lifted transform.
4.3. MGARD
@@
MGARD (Ainsworth et al., 2018, 2019a, 2019b, 2020; Gong et al., 2023) is a multilevel data compressor based on finite element analysis and wavelet theories. It treats the data as a piecewise multilinear function defined on the input data grid and iteratively decomposes the data into coarse representations in a set of hierarchical grids. The decomposition procedure is as follows. Starting with the original data and input grid, MGARD will compute the piecewise linear interpolation using data from the lower-level grid and then subtract the interpolation values from current data to obtain multilevel coefficients. These coefficients are then projected to the lower-level grid to compute correction, which roughly approximates the loss of missing nodes using the lower-level grid. The correction then is added to data in the lower-level grid to form the lower-level representation. This process is repeated until the lowest level is reached. All the multilevel coefficients are then fed to a Huffman encoder and a lossless encoder for size reduction.
MGARD can be applied to uniform/nonuniform structured and unstructured grids (Ainsworth et al., 2020) because of the general data decomposition theory. In addition to providing general error controls (such as absolute error and $L^{2}$ error) on raw data, MGARD features error controls on derived quantities such as bounded-linear analysis (Ainsworth et al., 2019b). It also provides error control for more complex derived quantities using a postprocessing method (Lee et al., 2022; Banerjee et al., 2022). The performance of MGARD is slightly slower than that of SZ and ZFP due to the higher computational complexity, but it provides portable GPU implementations across different vendors with high performance.
4.4. SPERR
@@
SPERR (Li et al., 2023) is a transform-based lossy compressor based on the CDF9/7 discrete wavelet transform (Cohen et al., 1992) and SPECK encoding algorithm (Pearlman et al., 2004), and it has both a pointwise error-bounding mode and a global quality thresholding mode. The compression pipeline of SPERR includes four stages: (1)) CDF9/7 wavelet transform; (2)) SPECK lossy encoding of wavelet coefficients; (3)) outlier encoding (only in error-bounding mode); and (4) zstd postprocessing of compressed data (optional). The decompression pipeline is an inverse of the compression pipeline with the decoding, detransform, and so on. The advantage of SPERR is that the hierarchical multidimension DWT in SPERR can effectively capture the relevance between data points, and it can often decorrelate the transformed coefficients to a great extent, which also brings a high compression ratio after the SPECK encoding. One limitation of SPERR is that the wavelet transform and the SPECK encoding processes have high computational costs,and hence its (sequential) execution speed is relatively low, typically around 30% of SZ3 (Zhao et al., 2021).
4.5. TTHRESH
@@
TTHRESH (Ballester-Ripoll et al., 2019) is a lossy compressor that utilizes the Tucker decomposition, specifically higher-order singular value decomposition. Unlike other lossy HOSVD-based compressors (Suter et al., 2011, 2013; Ballester-Ripoll et al., 2015), which implement coarse-granularity slicewise truncation on the tensor core and factor matrices post-HOSVD, TTHRESH employs bit-plane coding across the entire set of HOSVD transform coefficients. This approach is complemented by run-length encoding (RLE) and arithmetic coding (AC). Notably, TTHRESH is capable of achieving significantly higher compression ratios compared with other compressors, especially for larger error bounds (i.e., higher compression ratios). This superior performance is largely attributable to HOSVD’s efficiency in capturing the global correlations within the dataset. However, TTHRESH exhibits much lower speed compared with other compressors due to the high computational complexity of HOSVD; for instance, it is $O(n^{4})$ for a 3D dataset with dimensions of $n^{3}$. Additionally, we note that TTHRESH does not offer pointwise error control; instead, it can control only the $l^{2}$ error (the sum of squared errors) because of the nature of HOSVD.
4.6. FPZIP
@@
FPZIP (Lindstrom et al., 2017) is an error-controlled lossy compressor developed based on the prediction-based compression model. It involves four steps: (1) It uses a Lorenzo predictor to predict the data value for each data point. (2) It computes the prediction residuals and maps these to integers. (3) After the mapping, a two-level compression scheme is applied on the residual integers. (4) It then applies a fast entropy coding (arithmetic coding) to improve the compression ratio. FPZIP does not support absolute error bound or decimal digit control, but it allows users to specify the number of bit planes (i.e., precision) to ignore, based on which the users can control the data distortion on demand. Specifically, when the precision is set to 32 for the single-precision floating-point dataset, all 32 bit-planes will be preserved, projecting essentially a lossless compression. The lower the precision value, the higher the data distortion and also the higher the compression ratio.
4.7. QoZ
QoZ (quality-oriented scientific compressor) (Liu et al., 2022b) is a derivation and upgrade of the SZ3 error-bounded lossy compressor. It focuses on improving the decompression data quality, as its name indicates, and it also supports compression autotuning according to user-specified quality metric targets. The technical outlines of QoZ are detailed as follows. First, addressing the decompression visualization quality issue of SZ3 caused by the inaccurate long-interval data prediction, QoZ merges losslessly stored anchor points into its data interpolator design to avoid long-interval data predictions. Second, to further improve the compression rate-distortion, QoZ introduces levelwise autotuning of interpolation configurations and error bounds by the user-specified quality metric target. Third, the most recent version of HPEZ (or QoZ 2.0) (Liu et al., 2024) brings several major updates to its data prediction design, including multidimensional interpolation, interpolation re-ordering, dimension autofolding, and blockwise interpolation tuning. Those flexible design modules make QoZ a highly adaptive scientific data compressor with different optimization levels exhibiting varying compression ratios and speeds In experiments, with the compression speed of around 60% to 100% of SZ3, QoZ can achieve around 50% to 300% compression ratio improvement over SZ3 under the same quality metric (such as PSNR) value.
4.8. FAZ
In past research works, several systematical evaluations have shown that different scientific lossy compressor archetypes have diverse advantages and limitations. For example, wavelet-based SPERR features extremely high compression ratios over quite a few scientific datasets but may present unsatisfactory compression ratios on certain data inputs; Interpolation-based SZ3 and QoZ have stable and decent compression ratios over all scientific datasets but cannot achieve as high compression ratios as SPERR does on SPERR’s well-performing datasets. To this end, FAZ (Liu et al., 2023c) is proposed to offer scientific data users one versatile data compressor, freeing them from the work of compressor evaluation and selection. FAZ features a hybrid design of compression framework and compression pipeline. With an integrated pipeline autotuning module; it can adaptively leverage the best-fit compression techniques for each separate input and autodetermine their corresponding parameters. Attributed by this design, it achieves state-of-the-art compression ratio and distortion among all existing scientific error-bounded lossy compressors.
4.9. SZx
SZx (Yu et al., 2022) features a novel design that composes only lightweight operations, such as bitwise operations, additions, and subtractions, and can support strict control of the compression errors within user-specified error bounds. Specifically, SZx splits the whole dataset into the fixed-length blocks (each with 128 elements) and goes over each block to check whether all the elements in the block can be represented by a single value (i.e., the mean of min value and max value in the block). If yes, this block is called a “constant” block, and the block of data would be represented/compressed by using this single value (a kind of filtering). Otherwise, the block is called a “non-constant” block, and a bit-truncation method (i.e., PDP) is applied to compress these data. SZx is an ultrafast error-bounded lossy compressor that can achieve 2–7$\times$ faster throughput compared with the second-best existing error-bounded lossy compressor while still reaching a high compression ratio.
4.10. AE-SZ
AE-SZ (Liu et al., 2021a) is one of the initial explorations into leveraging deep neural networks for error-bounded scientific lossy compression. Specifically, AE-SZ applies a convolutional sliced-Wasserstein autoencoder with generalized divisive normalization layers for the compression; and in AE-SZ the Lorenzo data predictor is also used as a supplement to the neural networks. In evaluations, AE-SZ outperforms several existing scientific lossy compressors with traditional techniques such as SZ2.1 (Liang et al., 2018b), ZFP (Lindstrom, 2014), and SZauto (Zhao et al., 2020b) in the settings of relatively large error bounds.
4.11. SRNN-SZ
To the best of our knowledge, SRNN-SZ (Liu et al., 2023b) is the first work of applying both super-resolution neural networks and transformers to scientific error-bounded lossy compression. As a prediction-based data compressor, SRNN-SZ has a data prediction scheme that is nearly the same as interpolation-based SZ3 and QoZ; however, the super-resolution neural networks can perform as the alternative for the interpolation within a single level (especially the last ones). SRNN-SZ trains a customized hybrid attention transformer on assorted scientific datasets and fine-tunes it on each scientific domain before the application of the network. Evaluations (Liu et al., 2023b) show that SRNN-SZ has achieved state-of-the-art compression rate-distortion on several low-compressibility scientific datasets.
4.12. Digit Rounding and Bit Grooming
@@
Digit Rounding (dig, [n. d.]) and Bit Grooming (Zender, 2016) are two error-bounded lossy compressors, which both mainly adopt the bit-plane coding method. We describe these two compressors in detail as follows.
@@ 5. Customized Compressors for Specific Applications or Use Cases
5.1. Compression for Molecular Dynamics Simulations
Molecular dynamics (MD) simulations have become one of the most important research methods in many science domains, including physics, biology, and materials science. In biophysics and structural biology, MD simulations are commonly employed to study the behavior of macromolecules, such as proteins and nucleic acids, aiding in the interpretation of biophysical experiment results and the modeling of molecular interactions. In materials science, MD simulations enable researchers to model and predict the structural, thermal, and mechanical characteristics of materials at the atomic level, to help them understand phenomena such as material deformation, fracture mechanics, and phase transitions, offering insights that are often unattainable through direct experimental observation.
@@ 5.1. Compression for Molecular Dynamics Simulations
Lossy compression has been widely considered a promising solution to reduce the data volume of MD simulations. For example, GROMACS (Hess et al., 2008), which is one of the leading MD simulation packages, has had its lossy format XTC built-in for decades. However, designing lossy compressors for MD simulations also presents unique challenges. First, the dominant type of data that needs to be stored in MD simulations is the particle trajectory, which is made of multiple frames (snapshots) of particle coordinates in a 3D space. Other data, including the particle velocities, forces, and system topology, either are not required in many cases or take much less storage than the trajectory. Compared with the structured mesh (regular multidimensional grid) that many other scientific applications use, the trajectory format is different: it stores discrete coordinates whereas the mesh stores continuous values. As such, the trajectory format is not well studied in terms of compression—most of the leading lossy compressors, including SZ (Zhao et al., 2021) and ZFP (Lindstrom, 2014), focus on the continuous values from the structured mesh. Second, although the MD trajectory has a temporal dimension, it is not feasible to treat the data as time series and compress accordingly. One reason is that random access (access data from randomly selected frames) is usually required for postanalysis, such that the compression needs to be done in batches, each containing only a limited number of frames. Another reason is that the frames may be saved in irregular (random) intervals; therefore the data from MD simulations may not be as continuous as in regular time series.
@@
5.2. Compression for Quantum Chemistry Simulations
Quantum chemistry applications may produce extremely large amounts of data (such as petabytes of data (Gok et al., 2018)) during execution on parallel systems. General Atomic and Molecular Electronic Structure System (GAMESS) (gam, 2020) is a typical example. In GAMESS, the Schrödinger differential equation needs to be solved to obtain the wavefunction that contains all the information about a chemical system. The most expensive step in this procedure involves computation of two-electron repulsion integrals (ERIs), which takes about 87% time of Hartree–Fock computation time in GAMESS (Gok et al., 2018). This step also projects a high storage requirement because it scales as $O$($N^{4}$) with the size of the chemical system. ERIs are required by each time step during the simulation, but they cannot always be kept in memory because of limited memory capacity, so they need to be recomputed from scratch at every iteration.

### v2

@@ 5. General-Purpose Lossy Compressors for Scientific Data
SZ-family compressors. SZ (Di and Cappello, 2016; Tao et al., 2017c; Zhao et al., 2021; Liang et al., 2021a; Di and Cappello, 2018) is a prediction-based error-bounded lossy compressor. In fact, it is not only a compression library/software but also a flexible composable framework allowing users to customize specific compression pipelines according to their datasets or use cases. SZ’s compression pipeline generally is composed of four stages: pointwise data prediction, quantization, variable-length encoding and lossless encoding. For different domain datasets and use cases, the SZ developers have developed many predictors, including Lorenzo, linear regression, dynamic spline interpolation (Zhao et al., 2021), and also using ZFP transform as a predictor (Liang et al., 2019a). In addition to the data prediction stage, SZ developers explored the possibilities of improving compression capability by other stages, for example, different quantization methods such that the users can set various error bounds for multiple value ranges in one dataset (Liu et al., 2021b; Liu and et al., 2021; Liu et al., 2022a). SZ adopts Huffman+Zstd (Di and Cappello, 2016; Tao et al., 2017c; Liang et al., 2018b) to compress the quantization bins because this method projects the best trade-off between the compression ratio and compression speed (Liu et al., 2021c). SZ also supports pointwise relative error bound compression (Liang et al., 2018a), by leveraging a preprocessing step to transform the original data domain to the logarithm domain and then executing absolute error-bounded compression on top of it. To accelerate the throughput, an efficient fusion of logarithm was developed in (Zou et al., 2019, 2020).
Based on SZ3 framework, the SZ team has developed many variants to adapt to diverse user-demands and use-cases. For example, QoZ (quality-oriented compressor) (Liu et al., 2022b) supports compression autotuning according to user-specified quality metric targets. QoZ2.0/HPDZ (Liu et al., 2024) brings several major updates to its data prediction design, including multidimensional interpolation, interpolation re-ordering, dimension autofolding, and blockwise interpolation tuning, which can achieve 50% to 300% compression ratio improvement under the same quality metric (such as PSNR). FAZ (Liu et al., 2023c) combines SZ3 model with wavelet transform and can also adaptively leverage the best-fit compression techniques for each separate input and autodetermine their corresponding parameters. FAZ features extremely high compression ratios over SZ3 ( at a cost of compression throughput): it can get 3-4$\times$ compression ratio as high as that of QoZ on the SEGSalt dataset as shown in (Liu et al., 2023c). In addition, the SZ team also explored the possibility of leveraging DNN to improve compression ratios under the SZ3 framework: e.g., AE-SZ (Liu et al., 2021a) leverages AutoEncoder and SRNN-SZ (Liu et al., 2023b) adopts super-resolution technique. In addition, the SZ team also developed several ultra-fast versions for both CPU and GPU devices. SZx (Yu et al., 2022), for example, composes only lightweight operations, such as bitwise operations, additions, and subtractions, so that it can get a very high compression/decompression speed (about 4$\times$ as fast as SZ2 and ZFP in general). CuSZp (Huang et al., 2023a, 2024a) can get a very high end-to-end performance on GPU for both compression and decompression (200-400GB/s on A100 in general), still maintaining a relatively high compression ratio.
@@
ZFP-family compressors. ZFP (Lindstrom, 2014) is a transform-based error-bounded lossy compressor, which supports two error control methods: fixed-accuracy (i.e., absolute error bound) and fixed precision. ZFP splits the whole dataset into many fixed-size blocks (e.g., 4$\times$4$\times$4 for a 3D dataset) and then executes three steps in each block: (1) preprocessing (DT): align the values in a block to a common exponent and convert the floating-point values to a fixed-point representation; (2) (near)orthogonal block transform (OWT): use orthogonal transform to decorrelate data; and (3) embedded coding (BPC): order and encode the transform coefficients by the embedded coding. To achieve the best trade-off between decorrelation efficiency and speed, developers of ZFP explored multiple transforms using a parametric description and identified a near-orthogonal one to use in practice. Their embedded encoding is a variation of BPC, where the coefficients are divided into separate groups based on their locations and then encoded in the granularity of a group. In general, ZFP features high throughput both both compression and decompression because of the performance optimization strategies in its implementation, such as lifted transform. ZFP also features multiple variants to suit different devices or use-cases: e.g., ZFP-V and DE-ZFP are two hardware-optimized lossy compressor developed based on ZFP design principle; ZFP-X is an accelerated version with remodeled embedded coding.
@@
MGARD-family compressors. MGARD (Ainsworth et al., 2018, 2019a, 2019b, 2020; Gong et al., 2023) is a multilevel data compressor based on finite element analysis and wavelet theories. It treats the data as a piecewise multilinear function defined on the input data grid and iteratively decomposes the data into coarse representations in a set of hierarchical grids. The decomposition procedure is as follows. Starting with the original data and input grid, MGARD will compute the piecewise linear interpolation using data from the lower-level grid and then subtract the interpolation values from current data to obtain multilevel coefficients. These coefficients are then projected to the lower-level grid to compute correction, which roughly approximates the loss of missing nodes using the lower-level grid. The correction then is added to data in the lower-level grid to form the lower-level representation. This process is repeated until the lowest level is reached. All the multilevel coefficients are then fed to a Huffman encoder and a lossless encoder for size reduction. MGARD can be applied to uniform/nonuniform structured and unstructured grids (Ainsworth et al., 2020) because of the general data decomposition theory. In addition to providing general error controls (such as absolute error and $L^{2}$ error) on raw data, MGARD features error controls on derived quantities such as bounded-linear analysis (Ainsworth et al., 2019b), and also non-linear functions in MGARD-Lambda version (mga, [n. d.]). It also provides error control for more complex derived quantities using a postprocessing method (Lee et al., 2022; Banerjee et al., 2022). The performance of MGARD is slightly slower than that of SZ and ZFP due to the higher computational complexity, but it provides an accelerated version for structured dataset (MGARD+ version (Liang et al., 2021b) and a portable GPU implementations across different vendors with high performance.
@@
SPERR. SPERR (Li et al., 2023) is a transform-based lossy compressor based on the CDF9/7 discrete wavelet transform (Cohen et al., 1992) and SPECK encoding algorithm (Pearlman et al., 2004), and it has both a pointwise error-bounding mode and a global quality thresholding mode. The compression pipeline of SPERR includes four stages: (1)) CDF9/7 wavelet transform; (2)) SPECK lossy encoding of wavelet coefficients; (3)) outlier encoding (only in error-bounding mode); and (4) zstd postprocessing of compressed data (optional). The decompression pipeline is an inverse of the compression pipeline with the decoding, detransform, and so on. The advantage of SPERR is that the hierarchical multidimension DWT in SPERR can effectively capture the relevance between data points, and it can often decorrelate the transformed coefficients to a great extent, which also brings a high compression ratio after the SPECK encoding. One limitation of SPERR is that the wavelet transform and the SPECK encoding processes have high computational costs,and hence its (sequential) execution speed is relatively low, typically around 30% of SZ3 (Zhao et al., 2021).
@@
TTHRESH. TTHRESH (Ballester-Ripoll et al., 2019) is a lossy compressor that utilizes the Tucker decomposition, specifically higher-order singular value decomposition. Unlike other lossy HOSVD-based compressors (Suter et al., 2011, 2013; Ballester-Ripoll et al., 2015), which implement coarse-granularity slicewise truncation on the tensor core and factor matrices post-HOSVD, TTHRESH employs bit-plane coding across the entire set of HOSVD transform coefficients. This approach is complemented by run-length encoding (RLE) and arithmetic coding (AC). Notably, TTHRESH is capable of achieving significantly higher compression ratios compared with other compressors, especially for larger error bounds (i.e., higher compression ratios). This superior performance is largely attributable to HOSVD’s efficiency in capturing the global correlations within the dataset. However, TTHRESH exhibits much lower speed compared with other compressors due to the high computational complexity of HOSVD; for instance, it is $O(n^{4})$ for a 3D dataset with dimensions of $n^{3}$. Additionally, we note that TTHRESH does not offer pointwise error control; instead, it can control only the $l^{2}$ error (the sum of squared errors) because of the nature of HOSVD.
@@
FPZIP. FPZIP (Lindstrom et al., 2017) is an error-controlled lossy compressor developed based on the prediction-based compression model. It involves four steps: (1) It uses a Lorenzo predictor to predict the data value for each data point. (2) It computes the prediction residuals and maps these to integers. (3) After the mapping, a two-level compression scheme is applied on the residual integers. (4) It then applies a fast entropy coding (arithmetic coding) to improve the compression ratio. FPZIP does not support absolute error bound or decimal digit control, but it allows users to specify the number of bit planes (i.e., precision) to ignore, based on which the users can control the data distortion on demand. Specifically, when the precision is set to 32 for the single-precision floating-point dataset, all 32 bit-planes will be preserved, projecting essentially a lossless compression. The lower the precision value, the higher the data distortion and also the higher the compression ratio.
HAE. The Hierarchical AE-based Lossy Compressor (HAE) (Le et al., 2024) is a representative compressor designed using an Auto-Encoder (AE) to handle scientific datasets. As noted earlier, modern lossy compressors often integrate multiple compression techniques to achieve high compression ratios and quality, and HAE is no exception. It consists of four main steps: (1) data standardization, (2) encoding and decoding using a hierarchical Auto-Encoder, (3) quantization of the latent vector, and (4) error-thresholding of the reconstructed data. The data standardization step serves as a form of “domain transformation,” converting each original data point into a standardized value as a preprocessing step for compression. The hierarchical Auto-Encoder reduces the data dimensionality iteratively, followed by quantization. Finally, the error-thresholding step identifies “outlier” data points that exceed user’s error bound and preserves their original values to maintain accuracy.
@@
Digit Rounding and Bit Grooming. Digit Rounding (dig, [n. d.]) and Bit Grooming (Zender, 2016) are two error-bounded lossy compressors, which both mainly adopt the bit-plane coding method. We describe these two compressors in detail as follows.
@@ 6. Customized Compressors for Specific Applications or Use Cases
6.1 Compression for Molecular Dynamics Simulations
Molecular dynamics (MD) simulations have become one of the most important research methods in many science domains, including physics, biology, and materials science.
@@ 6.1 Compression for Molecular Dynamics Simulations
Lossy compression has been widely considered a promising solution to reduce the data volume of MD simulations. For example, GROMACS (Hess et al., 2008), which is one of the leading MD simulation packages, has had its lossy format XTC built-in for decades. However, designing lossy compressors for MD simulations also presents unique challenges. First, the dominant type of data that needs to be stored in MD simulations is the particle trajectory, which is made of multiple frames (snapshots) of particle coordinates in a 3D space. Other data, including the particle velocities, forces, and system topology, either are not required in many cases or take much less storage than the trajectory. Compared with the structured mesh (regular multidimensional grid) that many other scientific applications use, the trajectory format is different: it stores discrete coordinates whereas the mesh stores continuous values, so that the existing error-bounded lossy compressors such as SZ and ZFP may not work effectively. Second, although the MD trajectory has a temporal dimension, it is not feasible to treat the data as time series to compress because of the requirement of random access (access data from randomly selected frames) for postanalysis, such that the compression needs to be done in batches, each containing only a limited number of frames.
@@
6.2 Compression for Quantum Chemistry Simulations
Quantum chemistry applications may produce extremely large amounts of data (such as petabytes of data (Gok et al., 2018)) during execution on parallel systems. General Atomic and Molecular Electronic Structure System (GAMESS) (gam, [n. d.]) is a typical example. In GAMESS, the Schrödinger differential equation needs to be solved to obtain the wavefunction that contains all the information about a chemical system. The most expensive step in this procedure involves computation of two-electron repulsion integrals (ERIs), which takes about 87% time of Hartree–Fock computation time in GAMESS (Gok et al., 2018). This step also projects a high storage requirement because it scales as $O$($N^{4}$) with the size of the chemical system. ERIs are required by each time step during the simulation, but they cannot always be kept in memory because of limited memory capacity, so they need to be recomputed from scratch at every iteration.